for h(x) = 2/3 x -5 find the value of x for which h(x) = -7


1
Expert's answer
2022-03-18T09:23:48-0400

Solution


Given that


"h\left( x \right) = \frac{2}{3}x - 5\"


Now for "h\left( x \right) = - 7\"


We have


"\begin{gathered}
 h\left( x \right) = \frac{2}{3}x - 5 \\
 \Rightarrow & - 7 = \frac{2}{3}x - 5 \\
 \Rightarrow & - 7 \times 3 = \frac{2}{3}x \times 3 - 5 \times 3 \\
 \Rightarrow & - 21 = 2x - 15 \\
 \Rightarrow & - 21 + 15 = 2x \\
 \Rightarrow & - 6 = 2x \\
 \Rightarrow & x = - 3 \\ 
\end{gathered} \"


Hence the value of "x" for which "h(x)=-7" is "x=-3"


We can write it as


"h(-3)=-7"